Capital asset pricing model (CAPM) For the asset s
Answer #1

Where: Risk free rate Pa Beta of the security mExpected market return

Required return = 0.03 + 1.8 ( 0.11 - 0.03)

= 0.03 + 1.8 * 0.08

= 0.03 +0.144

= 17.4 %